Grade 12Physical Chemistry

An element crystallized in a structure having a fccunit cell of an edge length200pm. Calculate its density if 400g of this element contain 48into10power23 atoms

Samuel Ashish Colvin
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
Density of unit cell is given as:
d = z x M /a^3 x Na
    z = (6*1/2 + 8/8) => 4
Therfore,
d = 4*400*Na/48*10^23*Na*200*10^-12
   = 1600/48*10^11*200
   = 8/48*10^11
   = 10^-11/6 g/m^3